polynomial function of least degree with integral coefficients that has the given zeros -5+i
Answer by ikleyn(52824)

.
(x-(-5+i))*(x-(-5-i)) = ((x+5)-i)*(x+5)+i) = (x+5)^2 + 1 = x^2 + 10x + 26.